OpenStudy (lena772):

Choose the equation below that represents the line passing through the point (-3, -1) with a slope of 4. y - 1 = 4(x - 3) y + 1 = 4(x + 3) y - 3 = 4(x + 1) y + 3 = 4(x - 1)

hartnn (hartnn):

you can do this

hartnn (hartnn):

you know the point slope form, right ?

OpenStudy (lena772):

I know but I always screw up like last time.

hartnn (hartnn):

lol, that was once

OpenStudy (lena772):

y-y1=m(x-x1)

hartnn (hartnn):

yes

OpenStudy (lena772):

y--1=4(x--3) y+1=4(x+3)

OpenStudy (lena772):

Is that a solution? I'm afraid to look lol

hartnn (hartnn):

\(\huge \color{red } \checkmark \)